OpenStudy (anonymous):

Determine the area under the standard normal curve that lies to the left of the following z score. Use a table if necessary. Round your answer to four decimal places. z = 1.57

OpenStudy (anonymous):

Do you have a standard normal table? This can also be calculated precisely using a calculator that has statistics functions.

OpenStudy (anonymous):

There is a calculator available here: http://www.danielsoper.com/statcalc3/calc.aspx?id=2 It gave a value correct to 7 decimal places. My TI-84 computes it out to 10 places. Most statistical tables will only provide 4 or 5 decimal digits of accuracy, which is usually sufficient.

OpenStudy (anonymous):

You could also evaluate the following integral: \[\frac{1}{2}+\frac{1}{\sqrt{\pi}}\int\limits_{0}^{z/2}e^{-z^2}dz : z=1.57\]